OpenStudy (anonymous):

At the city museum, child admission is $5.80 and adult admission is $9.10 . On Wednesday, 155 tickets were sold for a total sales of $1083.80 . How many child tickets were sold that day?

OpenStudy (igreen):

We can make a system of equations. 5.80x + 9.10y = 1083.80 x + y = 155 Where 'x' is the number of child tickets and 'y' is the number of adult tickets.

OpenStudy (igreen):

x + y = 155 Subtract 'x' to both sides: y = -x + 155 Plug in '-x + 155' for 'y' in the first equation: 5.80x + 9.10y = 1083.80 5.80x + 9.10(-x + 155) = 1083.80 Simplify & solve for 'x'

OpenStudy (anonymous):

99?

OpenStudy (igreen):

Correct!

OpenStudy (igreen):

x = 99 And we said 'x' represents the amount of child tickets.

OpenStudy (anonymous):

thank you